OpenStudy (anonymous):

help please 1. Find the distance between the points (-4, 6) and (8, -12).Round your solution to 2 decimal points.

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

yeah

hartnn (hartnn):

so plug in values and tell me what u get ?

hartnn (hartnn):

do u need help with simplification ?

OpenStudy (anonymous):

16 n 36

hartnn (hartnn):

whats that ? and NO

OpenStudy (anonymous):

distance= (-4+8)^2 (6-12)^2 (4)^2 (-6)^2 (16) (36)

hartnn (hartnn):

there is a -*minus* sign in between -4-(8) = -12 (-12)^2 = ?

OpenStudy (anonymous):

+144

hartnn (hartnn):

6-(-12) = ?

OpenStudy (anonymous):

-6?

hartnn (hartnn):

6-(-12) = 6+12 = ?

OpenStudy (anonymous):

oh 18

hartnn (hartnn):

now square it

OpenStudy (anonymous):

but how do you write this out then? i got it wrong distance= (-4-(8) )^2 (6-12)^2 (-12)^2 (-6)^2 (144) (36)

hartnn (hartnn):

its 18^2 \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2} \\ \huge \huge d=\sqrt{(12)^2+(18)^2}=?\)

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

144+ 324

hartnn (hartnn):

yes, then take square root

hartnn (hartnn):

what u get ?

hartnn (hartnn):

did u get 21.63 ?

OpenStudy (anonymous):

no?

hartnn (hartnn):

whats 144+324 = ?

hartnn (hartnn):

its 468 and square root of 468 is 21.63 using calculator......

OpenStudy (anonymous):

isnt it 470?

hartnn (hartnn):

no, 144+324=468